OpenStudy (anonymous):

Julius has a garden that has dimensions of 12 ft by 20 ft. Julius needs more room to plant some peppers, and determines the garden will need a new area of 360 square feet. How much should he increase the length and width by, if he wants to increase them by the same amount? Round your answer to the nearest tenth. A. 3.8 feet B. 3.6 feet C. 3.4 feet D. 35.4 feet

OpenStudy (anonymous):

Well you can determine the current area with 12(20) = 240 Then you can add the choices to the length and width to calculate 360

OpenStudy (anonymous):

For example, let's try option A (12+3.8)(20+3.8) = ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its not A.

OpenStudy (cwrw238):

KIrbyLegs method is a good one

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

You can then try B and C now (12+3.6)(20+3.6) = ? (12+3.4)(20+3.4) = ? Thx cwrw. The only problem is that it is playing with the multiple choice answers

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

right kirby

OpenStudy (cwrw238):

another way to do it would be to solve the equation (12 + x)(20 + x) = 360 expand and rearrange to get 0 on RHS and solve using the quadratic formula